OpenStudy (anonymous):

A uniform solid disk of radius R and mass M is free to rotate on a frictionless pivot through a point on its rim. Find the speed of its center of mass.

OpenStudy (anonymous):

ugh so I = Icm + Iaxis = 1/2MR^2 + MR^2 = 3MR^2/2

OpenStudy (ghazi):

how can you calculate speed, you need angular velocity for that and use V=R*W also you don't have angular acceleration or displacement, you just can calculate center of mass but not the speed of rotation

OpenStudy (anonymous):

yea.. i think you are supposed to calculate the moment of inertia!